# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ IV-E Takeaways
Conforming to the usage trends of users, the collaboration of multiple agents has become a clear development direction. In this context, agent communication becomes the foundation of future AI ecosystem. Based on the communication entity, we classify agent communication into three different types: user-agent interaction, agent-agent communication, and agent-environment communication, and use an example to illustrate its entire lifecycle. This classification can naturally distinguish communication with similar vulnerability characteristics, providing structured research paradigm for future studies.
Figure 5: Taxonomy of our survey of agent communication protocols, security risks, and defense countermeasures.
@@ V-A Protocols
PXP. PXP protocol [242] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields. It is worth to mention that PXP is not customized for LLM-driven agents, but we think its design has inspirational meaning for agent communication. Therefore, we finally discuss it in this paper. PXP deploys a “two-way intelligibility” mechanism as its core and uses four message tags, namely RATIFY, REFUTE, REVISE, and REJECT, to regulate the interaction between human experts and agent. At the beginning of the interaction, the agent initiates a prediction and provides an explanation first. Subsequently, the two parties communicate alternately. A finite-state machine is used to calculate the message tags and update the context based on the prediction matching (MATCH) and explanation agreement (AGREE) situations. PXP uses a black board system to store data, messages, and context information. The process continues until the message limit is reached or specific termination conditions occur. The effectiveness of PXP has been verified in the scenarios of radiology and drug discovery.
@@
Spatial Population Protocols. The Spatial Population Protocols is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems. Similar to PXP, strictly speaking, this work is not designed for LLM-driven agent systems. However, since it may benefit agents requiring location services, we also discuss it in this paper. Spatial Population Protocols allows agents to obtain pairwise distances or relative position vectors when interacting in Euclidean space. Each agent can store a fixed number of coordinates. During interaction, in addition to exchanging knowledge, geometric queries can also be performed. Through the multi-contact epidemic mechanism, leader election, and self-stabilizing design, it enables n anonymous robots to achieve efficient localization from their respective inconsistent coordinate systems to a unified coordinate consensus, providing a scalable framework for robot collaboration in dynamic environments.
@@
AG-UI. AG-UI [206] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism. The front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols). The client invokes the RUN interface of the protocol layer to sends requests to the agent. When the agent processes the request, it generates a streaming event and returns it to the AG-UI client. Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events. The AG-UI client handles different types of responses by subscribing to the event stream. Agents can transfer context between each other to maintain the continuity of the conversation. All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication.
@@
Besides, please note that the previous survey [299] regards CrowdES [20] as an agent-related protocol. After our careful study, we think CrowdES is a framework for generating and evaluating simulated crowds and real-world crowds, which is not suitable to be discussed in the field of agent communication. Therefore, we will not list it in this paper.
V-B Security Risk Analysis
@@
TABLE IV: Mapping of User-Agent Communication Risks to Their Characteristics and Defense Strategies
Risk Type Representative Threats Attack Characteristics Defense Strategy
Text-Based Attacks Direct Prompt Injection Overrides agent behavior through explicit adversarial inputs Input filtering (intent detection, perplexity scoring, safety classifiers)
Indirect Prompt Injection Stealthy injection via external sources (RAG, web metadata) Source vetting, risk scoring (e.g., TrustRAG), sandboxing
@@
Jailbreak Prompts Bypasses safety via role-play, obfuscation, multi-turn tricks Output filtering (e.g., GPT-Fuzzer, ShieldLM)
Cross-Agent Prompt Propagation Malicious prompts propagate across agents in MAS Message isolation, sanitization agents, structured content filtering
@@
Multimodal Attacks Image-Based Attacks Embeds adversarial cues in visuals while text remains benign Image purification (e.g., cropping, compression, diffusion)
@@
Audio-Based Attacks Injects commands via adversarial waveforms or speech mimicry Audio sanitization (e.g., noise, resampling, filtering)
Cross-Modal Inconsistency Exploits mismatch between modalities (e.g., image vs. text) Embedding alignment checks, OCR + caption validation
Privacy Leakage Identity Inference Infers identity via face, gesture, or voiceprint recognition Data minimization, anonymization (e.g., IdentityDP)
Prompt-Induced Disclosure Triggers agent to leak sensitive content via crafted prompts Prompt-level leakage detection (e.g., GenTel-Shield)
Behavioral Profiling Tracks patterns across modalities to infer user behavior Feature disentanglement, modality isolation, masking
Denial of Service (DoS) Token Flooding Induces long repetitive output to exhaust system resources Quota control, length prediction, truncation mechanisms
OverThinking Attacks Forces agent into unnecessary reasoning, increases latency Depth-limiting, adversarial training, anomaly detection
@@
According to our analysis, the user-agent interaction shows significant multimodal characteristic, as users often interact through text, images, and videos. Consequently, the security risks in user-agent interaction primarily stem from these insecure inputs.
V-B1 Text-Based Attacks.
@@
  • Prompt Injection refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources. It can be classified into two categories: Direct Prompt Injection and Indirect Prompt Injection. Direct prompt injection refers to user input that explicitly alters agent’s behavior in unintended ways. Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [171, 170, 236, 210, 173] to override the original prompt and subvert agent’s intended behavior. In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [83, 42]. For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [41, 348, 13, 28, 50]; in web-augmented agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agent’s response [65, 37].
@@
V-B3 Privacy Leakage.
While multimodal systems improve user experience, they collect rich sensory data containing identity, emotion, and behavior patterns. Without effective data governance, attackers may exploit visual tracking, gesture recognition, or cross-modal inferences to reconstruct identities or infer psychological states, enabling passive profiling or behavioral prediction [260, 334, 199, 338, 154, 118]. Want et al. [264] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems. MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture. By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure. Li et al. [152] reveal that commercial agents are vulnerable to simple yet dangerous attacks. These attacks manipulate malicious content (such as fake product pages, forged academic papers) on trusted platforms to induce agents to leak users’ credit card information, download viruses, send phishing emails, and even mistakenly synthesize toxic substances like nerve gas.
V-B4 Denial of Service.
@@
V-C Defense Countermeasure Prospect
@@
We will investigate the possible defense measures that can be taken to address the security risks in the user-agent interaction.
V-C1 Countermeasures for Text-Based Attacks
@@
Input and Output Filtering. Before user inputs are processed by the agent system, multiple approaches can be conducted for semantic-level input safety review. For example, methods based on intent analysis [330, 274], perplexity calculation [116], and fine-tuned safety classifiers [157, 333, 314, 112] can be employed to identify attack instructions and malicious intentions in the input stage. After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [333, 314, 112, 190, 305], to ensure alignment with safety objectives.
@@
V-C2 Countermeasures for Multimodal Attacks
@@
Input Pre-processing and Sanitization. As the first line of defense, input sanitization aims to clean potentially malicious content before it is processed by the model.
  • Image Purification: To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals. These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [110, 47, 291]. Although lightweight, such operations can significantly degrade pixel-level adversarial patterns meticulously crafted by attackers, thereby reducing the attack success rate. In addition, diffusion models can be used to reconstruct the input image, effectively ”washing out” subtle and imperceptible adversarial perturbations [198].
@@
V-C3 Countermeasures for Privacy Leakage
@@
Data Minimization and Anonymization. During the multimodal data collection phase, a strict data minimization principle should be enforced, ensuring that only the information necessary for task completion is collected. Sensitive biometric data (e.g., facial features, voiceprints, gesture patterns) should be processed using differential privacy or k-anonymity techniques to mitigate the risk of identity reconstruction. Besides, a hierarchical data access control mechanism should be established to ensure that each system component can access only the minimal dataset required for its functionality. To protect sensitive biometric features such as facial information, Wen et al. [282] proposes a differential privacy-based anonymization framework IdentityDP to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
@@
Cross-modal Inference Restriction. To mitigate the risks of identity inference through cross-modal correlations, it is essential to design modality-level information isolation mechanisms. This can be achieved by introducing noise perturbations or feature disentanglement techniques to disrupt the direct associations between different modalities while preserving overall system functionality. In addition, dynamic feature masking can be implemented by periodically altering data representations, thereby increasing the difficulty for adversaries to perform long-term behavioral analysis.
V-C4 Countermeasures For DoS
@@
V-D Takeaways
User-agent interaction enables agents to process multimodal inputs, such as text, images, audio, and the combinations of them. Since this process needs to directly face the uncertain inputs from diverse users, its security risks are severe. We divide the existing risks into text-based attacks, multimodal attacks, privacy leakage, and denial-of-service. Then we detailedly outlook on the defenses countermeasures for each risk. Overall, this interaction process is critical for bridging human intent with agent execution, but its security requires long-term study.

### v2

@@ IV-E Takeaways
Conforming to the usage trends of users, the collaboration of multiple agents has become a clear development direction. In this context, agent communication becomes the foundation of the future AI ecosystem. Based on the communication entity, we classify agent communication into three different types: user-agent interaction, agent-agent communication, and agent-environment communication. This classification can naturally distinguish communication with similar vulnerability characteristics, providing a structured research paradigm for future studies.
@@ V-A Protocols
PXP. PXP protocol [256] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields. It is worth mentioning that PXP is not customized for LLM-driven agents, but we think its design has inspirational meaning for agent communication. Therefore, we finally discuss it in this paper. PXP deploys a “two-way intelligibility” mechanism as its core and uses four message tags, namely RATIFY, REFUTE, REVISE, and REJECT, to regulate the interaction between human experts and agents. At the beginning of the interaction, the agent initiates a prediction and provides an explanation first. Subsequently, the two parties communicate alternately. A finite-state machine is used to calculate the message tags and update the context based on the prediction matching (MATCH) and explanation agreement (AGREE) situations. PXP uses a blackboard system to store data, messages, and context information. The process continues until the message limit is reached or specific termination conditions occur. The effectiveness of PXP has been verified in the scenarios of radiology and drug discovery.
@@
Spatial Population Protocols. The Spatial Population Protocols is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems. Similar to PXP, strictly speaking, this work is not designed for LLM-driven agent systems. However, since it may benefit agents requiring location services, we also discuss it in this paper. Spatial Population Protocols allow agents to obtain pairwise distances or relative position vectors when interacting in Euclidean space. Each agent can store a fixed number of coordinates. During interaction, in addition to exchanging knowledge, geometric queries can also be performed. Through the multi-contact epidemic mechanism, leader election, and self-stabilizing design, it enables n anonymous robots to achieve efficient localization from their respective inconsistent coordinate systems to a unified coordinate consensus, providing a scalable framework for robot collaboration in dynamic environments.
@@
AG-UI. AG-UI [217] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism. As shown in Figure 6, the front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols). The client invokes the RUN interface of the protocol layer to send requests to the agent. When the agent processes the request, it generates a streaming event and returns it to the AG-UI client. Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events. The AG-UI client handles different types of responses by subscribing to the event stream. Agents can transfer context between each other to maintain the continuity of the conversation. All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication.
Figure 6: The architecture of AG-UI.
@@
Besides, please note that the previous survey [318] regards CrowdES [24] as an agent-related protocol. After our careful study, we think CrowdES is a framework for generating and evaluating simulated crowds and real-world crowds, which is not suitable for discussion in the field of agent communication. Therefore, we will not list it in this paper.
V-B Security Risk Analysis: Malicious Users Against Benign Agents
@@
TABLE IV: The Security Risks in the User-Agent Interaction Phase and Their Characteristics
Stage Risk Type Representative Threats Attack Characteristics
Malicious users against benign agents Text-Based Attacks Prompt Injection Manipulate agent behavior by malicious prompts, not modifying model
@@
Jailbreak Prompts Bypass security constraints via role-play, obfuscation, multi-turn tricks
@@
Multimodal Attacks Image-Based Attacks Embed adversarial cues in visual inputs while text remains benign
@@
Audio-Based Attacks Inject commands via adversarial waveforms or speech mimicry
Privacy Leakage Inference Attacks Disclosure of agent information, such as topological structure
Denial of Service (DoS) Overthink Attacks Forces agent into unnecessary reasoning, repeat actions, or overthinking
Compromised agents against benign users Violation of User Privacy Exposure of Personal Information Expose a comprehensively aggregated user profile
Behavioral and Psychological Profiling Build detailed behavioral or psychological profiles from innocuous conversations
Psychological and Social Manipulation Belief and Opinion Shaping Manipulate user’s worldview, influence voting behavior, or radicalize beliefs
Social Engineering Leverage the knowledge of the user’s communication style, vocabulary, and relationships to conduct highly convincing impersonation attacks
Harmful Tasks Economic Manipulation Inflict subtle yet significant damage in professional or economic contexts
Behavioral and Psychological Profiling Build unauthorized profiles and infer highly sensitive attributes
@@
According to our analysis, the inputs from users show a significant multimodal characteristic. For example, AG-UI explicitly illustrates its support for multimodal content, such as text, images, and videos. Consequently, we find that the security risks in user-agent interaction primarily stem from these insecure inputs.
V-B1 Text-Based Attacks
@@
  • Prompt Injection refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources. It can be classified into two categories: Direct Prompt Injection and Indirect Prompt Injection. Direct prompt injection refers to user input that explicitly alters the agent’s behavior in unintended ways. Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [182, 181, 249, 223, 184] to override the original prompt and subvert the agent’s intended behavior. In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [92, 46]. For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [45, 365, 14, 32, 54]; in web-augmented agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agent’s response [70, 41].
@@ V-B2 Multimodal Attacks
V-B3 Privacy Leakage
Without effective data governance, the rich sensory data may be exploited by malicious users to launch various forms of privacy breaches, posing significant risks to the confidentiality of the agent system. Want et al. [282] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems. MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture. By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure.
V-B4 Denial of Service
@@
V-C Security Risk Analysis: Compromised Agents Against Benign Users
Apart from the threats from malicious users, we point out that the risks caused by compromised agents against benign users are also worth attention. This section categorizes the destructive impact and security risks from compromised agents based on the user’s perspective.
V-C1 Violation of User Privacy
A compromised agent becomes a channel for data exfiltration, directly targeting the user’s sensitive information. The harm manifests in several ways:
  • Exposure of Personal Information: A compromised agent can be induced to leak the user’s Personally Identifiable Information (PII) that it has access to, such as the user’s name, email, address, and conversation history [292, 172, 273]. In more severe cases, this can extend to financial data like credit card numbers or passwords [12], leading to direct financial loss. What makes this threat particularly potent is the agent’s role as a central data aggregator. Since agents often integrate with multiple user data silos (e.g., email, calendar, cloud storage, and social media), a breach does not just expose isolated pieces of information. Instead, it allows for the exfiltration of a comprehensively aggregated user profile, where the potential for harm far exceeds the sum of its parts.
  • Behavioral and Psychological Profiling: A compromised agent can be manipulated to analyze the user’s inputs across sessions to build a detailed behavioral or psychological profile against their will. Moreover, a more insidious risk happens where the agent deduces highly sensitive attributes (e.g., health conditions, political affiliations, or undisclosed personal relationships) from seemingly innocuous conversational data that the user never explicitly provided [63, 91, 278, 274]. These disclosed profiles put users at the risk of manipulation, targeted scams, or social engineering.
V-C2 Psychological and Social Manipulation
Beyond simple data theft, a compromised agent can become a powerful tool for psychological manipulation, exploiting the user’s trust and the agent’s persuasive capabilities. This form of attack targets the user’s beliefs, decisions, and relationships.
  • Belief and Opinion Shaping: The agent can be instructed to subtly introduce biased information, conspiracy theories, or political propaganda into its responses over time. By personalizing the misinformation to the user’s psychological profile, the agent can effectively manipulate their worldview, influence their voting behavior, or radicalize their beliefs. This exploits the inherent persuasiveness of conversational AI. Park et al. [220] highlights how models can be used for manipulative purposes, including generating persuasive, deceptive content that is difficult for humans to detect. They note that AI can “simulate empathy” to build rapport before manipulating the user.
  • Sophisticated Social Engineering and Impersonation: A compromised agent has intimate knowledge of the user’s communication style, vocabulary, and relationships (gleaned from emails, messages, etc.). It can leverage this to conduct highly convincing impersonation attacks. For example, it could send a fraudulent email to the user’s colleague or family member, perfectly mimicking the user’s tone, to request a password reset, a fund transfer, or sensitive information. This attack is far more credible than generic phishing attempts. Greshake et al. [92] demonstrates how an agent can be poisoned by external data (like a webpage it summarizes) and then turned against the user or even used to attack other systems on the user’s behalf. This mechanism could be used to weaponize an agent for impersonation.
V-C3 Execution of Malicious and Harmful Tasks
Once compromised, an agent can be weaponized, transforming from a trusted assistant into an active executor of malicious tasks that can sabotage the user’s interests or directly endanger them, representing a significant risk escalation [192].
  • Economic Manipulation: The agent can be instructed to inflict subtle yet significant damage in professional or economic contexts. For a user relying on it for work, it could covertly introduce logical errors into computer code, provide flawed data in financial projections, or leak confidential business strategies discussed in conversations [221]. The harm is often latent and difficult to detect, potentially leading to professional failure or corporate espionage. This extends to broader market manipulation, where an agent could use the user’s social media accounts to automate large-scale disinformation campaigns, such as posting fake product reviews or spreading rumors to affect a company’s stock price, making the user an unwitting accomplice in a larger economic attack.
  • Malicious Guidance: A compromised agent can also be used as a direct vector for attacking the user’s digital environment. It can be triggered to generate scripts that download malware, trick the user into visiting phishing websites, or send highly convincing phishing emails on the user’s behalf, thereby damaging their reputation and spreading the attack to their contacts [273, 158]. In a more severe scenario, a jailbroken or manipulated agent can bypass its safety protocols to provide overtly harmful instructions. This includes generating tutorials for synthesizing toxic substances, creating malicious code on demand, or providing dangerously flawed medical or financial advice, directly threatening the user’s physical safety and financial stability [31, 183, 294, 174].
V-D Defense Countermeasure Prospect
@@
We will investigate the possible defense measures that can be taken to address the security risks in the user-agent interaction. Here we only present the defenses for attacks from malicious users; those for compromised agents (against users, other agents, and environments) are uniformly put in Section VI-C.
V-D1 Countermeasures for Text-Based Attacks
@@
  • Input and Output Filtering. Before user inputs are processed by the agent system, multiple approaches can be conducted for a semantic-level input safety review. For example, methods based on intent analysis [349, 293], perplexity calculation [126], and fine-tuned safety classifiers [166, 352, 333, 122] can be employed to identify attack instructions and malicious intentions in the input stage. After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [352, 333, 122, 202, 324], to ensure alignment with safety objectives.
@@
V-D2 Countermeasures for Multimodal Attacks
@@
  • Image Purification: To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals. These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [120, 51, 310]. Although lightweight, such operations can significantly degrade pixel-level adversarial patterns meticulously crafted by attackers, thereby reducing the attack success rate. In addition, diffusion models can be used to reconstruct the input image, effectively “washing out” subtle and imperceptible adversarial perturbations [210].
@@
V-D3 Countermeasures for Privacy Leakage
@@
  • Data Minimization and Anonymization. During the multimodal data collection phase, a strict data minimization principle should be enforced, ensuring that only the information necessary for task completion is collected. Sensitive biometric data (e.g., facial features, voiceprints, gesture patterns) should be processed using differential privacy or k-anonymity techniques to mitigate the risk of identity reconstruction. Besides, a hierarchical data access control mechanism should be established to ensure that each system component can access only the minimal dataset required for its functionality. To protect sensitive biometric features such as facial information, Wen et al. [301] propose a differential privacy-based anonymization framework, IdentityDP, to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
@@
  • Cross-modal Inference Restriction. To mitigate the risks of identity inference through cross-modal correlations, it is essential to design modality-level information isolation mechanisms. This can be achieved by introducing noise perturbations or feature disentanglement techniques to disrupt the direct associations between different modalities while preserving overall system functionality. In addition, dynamic feature masking can be implemented by periodically altering data representations, thereby increasing the difficulty for attackers to perform long-term behavioral analysis.
V-D4 Countermeasures For DoS
@@
V-E Takeaways
User-agent interaction enables agents to process multimodal inputs, such as text, images, audio, and their combinations. Since this process needs to directly face the uncertain inputs from diverse users, its security risks are severe. We divide the existing risks into risks from malicious users and risks from compromised agents. Then we detailedly discuss the defense countermeasures for risks from malicious users. Defenses for compromised agents will be introduced in the agent-agent communication section. Overall, user-agent interaction is critical for bridging human intent with agent execution, but its security requires long-term study.